OpenStudy (anonymous):

what is the square root of 6 divided by the square root of 5 minus square root of 3?

OpenStudy (anonymous):

Can you please create this expression using the equation tool? The wording is very ambiguous.

OpenStudy (anonymous):

\[\sqrt{6\div} \sqrt{5}-\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{6}\div \sqrt{5}-\sqrt{3} \] So what we do is multiply the top and bottom by \[\sqrt{5}+\sqrt{3} \] you will get \[\sqrt{6} (\sqrt{5}+\sqrt{3})/5 -3\] I think that simplifies to \[\sqrt{30}+\sqrt{18}/2\].... you can do the rest of the simplifying i am too lazy ~.~...